OpenStudy (anonymous):

Use the limit definition of the derivative: y = sqrt(3 x)+1

OpenStudy (anonymous):

Okay: The limit definition of a derivative is: \[\lim_{\Phi \rightarrow 0}\frac{f(x+\Phi)-f(x)}{\Phi}\] So: \[f(x+\Phi)=\sqrt{3x+3\Phi}\] Now you have: \[\lim_{\Phi \rightarrow 0}\frac{\sqrt{3x+3\Phi}-\sqrt{3x}}{\Phi}\] (From here on out there is an implied limit in the front (saves time typing it over and over). Multiply top and bottom by the conjugate of the top: \[\frac{(\sqrt{3x+3\Phi}-\sqrt{3x})(\sqrt{3x+3\Phi}+\sqrt{3x})}{\Phi(\sqrt{3x+3\Phi}+\sqrt{3x})}\] Giving: \[\frac{3x+3\Phi-3x}{\Phi(\sqrt{3x+3\Phi}+\sqrt{3x})}=\frac{3\Phi}{\Phi(\sqrt{3x+3\Phi}+\sqrt{3x})}=\frac{3}{\sqrt{3x+3\Phi}+\sqrt{3x}}\] Taking the limit we have: \[\lim_{\Phi \rightarrow 0}\frac{3}{\sqrt{3x+3\Phi}+\sqrt{3x}}=\frac{3}{\sqrt{3x}+\sqrt{3x}}=\frac{3}{2\sqrt{3x}}\]

OpenStudy (anonymous):

What happens to the 1?

OpenStudy (anonymous):

Well when you have: \[\sqrt{3x+3\Phi}+1-(\sqrt{3x}+1) \implies \sqrt{3x+3\Phi}+\sqrt{3x}\] The ones cancel out. I jumped ahead of myself in the algebra. Sorry.